OpenStudy (anonymous):

Express the following in terms of sinθ and/or cosθ. cosθ 1/secθ

zepdrix (zepdrix):

hmm dont remember your identities? c= what is secant in relationship to cosine? hmm

OpenStudy (anonymous):

i forgot

zepdrix (zepdrix):

\[\sec \theta=\frac{ 1 }{ \cos \theta }\] hmmmm so what would 1/sec = ? :o

OpenStudy (anonymous):

cos

OpenStudy (anonymous):

-.-

zepdrix (zepdrix):

\:D/

OpenStudy (anonymous):

LOL

OpenStudy (anonymous):

could have sworn i just answered this one maybe it was not clear

OpenStudy (anonymous):

No the first 2 were right thank youuu